# Practical Evaluation of Copula-based Survival Metrics: Beyond the Independent Censoring Assumption

### v2

@@ E.2 Integrated Brier Score
[…]
where $t_{\text{max}}$ is usually the maximum event time of the combined training and validation datasets. This metric is implemented in the SurvivalEVAL package [Qi et al., 2023b].
E.3 Mean Absolute Error
@@
MAE-Margin [Haider et al., 2020] estimates a “best guess” event time for censored subjects using the KM estimator [Kaplan and Meier, 1958]. For a subject censored at time $t_{i}$, the margin event time is computed as:
@@
This metric is implemented in SurvivalEVAL package [Qi et al., 2023b].
Appendix F Experimental Details
F.1 Data generation and preprocessing
@@
Synthetic: This paper utilizes a linear synthetic dataset with two competing events (i.e., event and censoring), both based on the parametric Weibull distribution as discussed in Foomani et al. [2023]. The datasets are characterized by the conditional hazard function and conditional survival probability of an event or censoring as follows:
$h_{E\mid\bm{X}}(t\mid\bm{X})$ $=\left(\frac{v}{\rho}\right)\left(\frac{t}{\rho}\right)^{v-1}\exp\left(g_{\Psi}(\bm{X})\right),$ (F.1)
$S_{E\mid\bm{X}}(t\mid\bm{X})$ $=\exp\left(-\left(\frac{t}{\rho}\right)^{v-1}\exp\left(g_{\Psi}(\bm{X})\right)\right),$ (F.2)
@@
where $v$ and $\rho$ represent the shape and scale parameters of the Weibull distribution, respectively. We use $v_{E}=4$, $v_{C}=6$, $\rho_{E}=17$, $\rho_{C}=19$. $g_{\Psi}(\bm{X})$ represents the risk function, which is a linear function in this work. The data generation process begins by initializing a copula with a predefined parameter $\theta$. We then sample $N$ sets of CDF observations from the copula, each containing two CDF probabilities (corresponding to each competing event, i.e., event or censoring). From these probabilities, we then calculate the corresponding times using the inverse function of (F.2). The data generation process is further described in Algorithm 2 in Foomani et al. [2023].
@@
METABRIC: The Molecular Taxonomy of Breast Cancer International Consortium (METABRIC) dataset [Curtis et al., 2012] contains survival information for breast cancer patients. This dataset includes a diverse range of feature sets that encompass clinical traits, expression profiles, copy number variation (CNV) profiles, and single nucleotide polymorphism (SNP) genotypes. All these features are derived from breast tumor samples collected during the METABRIC trial. The dataset can be downloaded from (https://www.cbioportal.org/study/summary?id=brca_metabric), and it does not have any missing values.
@@
MIMIC-IV: The Medical Information Mart for Intensive Care (MIMIC)-IV [Johnson et al., 2023] contains critical care data from patients admitted to hospital and intensive care units (ICU). We create a version using the MIMIC-IV database which contains patients that are alive at least 24 hours after being admitted to ICU. Their date of death is derived from hospital records or state records, which means, the cause of mortality is not limited to the reason for ICU admission. We follow the instructions in [Qi et al., 2023a] to process the dataset and the code is available in their GitHub repository: https://github.com/shi-ang/CensoredMAE.
@@
SEER: The SEER Program dataset [LAG et al., 2007] is a comprehensive collection of cancer patient data from approximately 49% of the U.S. population. Our study focuses on three subsets: SEER-brain, SEER-liver, and SEER-stomach, containing data on patients diagnosed with brain, liver, and stomach cancers, respectively. The goal is to model the time from diagnosis to failure events, such as death or disease progression. We preprocess the dataset as in [Qi et al., 2024]. Features analyzed include age, sex, cancer stage, grade, and various clinical factors. We excluded features with over 70% missing data, patients without follow-up times, duplicates, and those with a survival time of zero. The SEER cohort is available for download at www.seer.cancer.gov.
F.2 Survival learners
@@
CoxPH: The Cox Proportional Hazards model (CoxPH) is a semi-parametric survival analysis model [Cox, 1972], consisting of a non-parametric baseline hazard function $h_{0}(t)$ and a parametric partial hazard function $f(\bm{\theta},\bm{x}_{i})$, typically estimated using the Breslow method [Breslow, 1975]. The hazard function is $h(t\mid\bm{x}_{i})=h_{0}(t)\exp(f(\bm{\theta},\bm{x}_{i}))$, where $f$ is linear. The MLE of $\hat{\bm{\theta}}$ maximizes the Cox partial log-likelihood [Cox, 1972]. CoxPH is implemented from the scikit-survival package version 0.21.0.
GBSA: GBSA [Ridgeway, 1999] is a likelihood-based boosting model that optimizes an $\ell_{2}$-norm penalized Cox partial log-likelihood [Cox, 1972]. It updates candidate variables flexibly at each boosting step using an offset-based gradient boosting method, unlike traditional gradient boosting, which updates either a single component or all features for fitting the gradient. GBSA is implemented from the scikit-survival package version 0.21.0.
RSF: Random Survival Forests (RSF) [Ishwaran et al., 2008] extend decision trees for survival analysis by constructing multiple trees on bootstrapped samples. Each tree predicts an individual’s survival function, with splits based on criteria like the log-rank test. RSF handles censored data as right-censored and accounts for it during tree construction, with final predictions as the average of all trees’ survival functions. RSF is implemented from the scikit-survival package version 0.21.0.
DeepSurv: DeepSurv [Katzman et al., 2018] is a neural network implementation of the CoxPH model of the form $h\left(t\mid\bm{x}_{i}\right)=h_{0}\left(t\right)\exp\left(f\left(\bm{\theta},\bm{x}_{i}\right)\right)$, where $f\left(\bm{\theta},\bm{x}_{i}\right)$ denotes a risk score as a nonlinear function of the features, i.e.$f\left(\bm{\theta},\bm{x}_{i}\right)=\sigma(\bm{x}_{i}\bm{\theta})$, where $\sigma$ is a nonlinear activation function. Popular functions include the hyperbolic tangent (Tanh) function or the Rectified Linear Unit (ReLU). The MLE for $\hat{\bm{\theta}}$ is derived by numerically maximizing the partial Cox log-likelihood. DeepSurv is implemented from https://github.com/shi-ang/BNN-ISD/.
MTLR: Multi-Task Logistic Regression (MTLR) Yu et al. [2011] is a neural network implementation of a discrete-time model that predicts the survival distribution using a sequence of dependent logistic regressors. The number of discrete times is determined by the square root of the number of uncensored patients, and we use quantiles to divide those uncensored instances evenly into each time interval. MTLR is implemented from https://github.com/shi-ang/BNN-ISD/.
@@
Table 4: Hyperparameters for the survival learners.
F.3 Implementation Details
@@
Copula model. To validate our approach in semi-synthetic datasets, we train two copula models, Clayton and Frank, alongside two nonlinear Weibull proportional-hazards models for the marginal distributions $E$ and $C$, respectively. The Weibull models use a multilayer perceptron (MLP) as the backbone network architecture. We use a single layer with 32 hidden nodes and the ReLU activation function. For all experiments, we train the models for 10000 epochs using the Adam optimizer [Kingma and Ba, 2017] with a learning rate of 1e-3, batch normalization and a batch size of 1024. We use early stopping to terminate the training process if the validation loss does not improve for 100 consecutive epochs.
@@
Loss function. We estimate the model parameters using maximum likelihood estimation (MLE) under dependent censoring, as described in Foomani et al. [2023]. Given survival data $\mathcal{D}$, let $f_{T\mid\bm{X}}(\cdot)$ and $F_{T\mid\bm{X}}(\cdot)$ represent the conditional density and cumulative distribution functions, respectively, over the event horizon, with the event distribution $E$ and the censoring distribution ${C}$. Under independence censoring, the likelihood of the $i$-th instance experiencing the event is:
$\mathcal{L}(\mathcal{D})=\prod_{i=1}^{N}\left[\text{Pr}(E=t_{i},C\geq t_{i}\mid\bm{X}=\bm{x}_{i})^{\delta_{i}}\times\text{Pr}(E\geq t_{i},{C}=t_{i}\mid\bm{X}=\bm{x}_{i})^{1-\delta_{i}}\right]$ (F.3)
@@
However, when $E$ and $C$ are no longer independent, we can no longer rely on this clean decomposition of the likelihood. Instead, we introduce a copula $C(u_{1},u_{2})$ to Equation F.3 to adjust for the dependency between $E$ and $C$ by applying Lemma 2 from Foomani et al. [2023]. Now, as a function of $C(u_{1},u_{2})$ and taking the sum of logs, the (dependent) likelihood has the following form:
@@
$\mathcal{L}(\mathcal{D})=\sum_{i=1}^{N}\delta^{(i)}\;\text{log}\left[f_{{E}\mid\bm{X}}(t_{i}\mid\bm{x}_{i})\right]\;+\;\delta_{i}\;\text{log}\left[\frac{\partial}{\partial u_{1}}C(u_{1},u_{2})\rvert^{u_{1}=S_{E\mid\bm{X}}(t_{i}\mid\bm{x}_{i})}_{u_{2}=S_{C\mid\bm{X}}(t_{i}\rvert\bm{x}_{i})}\right]$ (F.4)
$+\;(1-\delta_{i})\;\text{log}\left[f_{{C}\mid\bm{X}}(t_{i}\mid\bm{x}_{i})\right]+(1-\delta_{i})\;\text{log}\left[\frac{\partial}{\partial u_{1}}C(u_{1},u_{2})\rvert^{u_{1}=S_{E\rvert\bm{X}}(t_{i}\rvert\bm{x}_{i})}_{u_{2}=S_{{C}\rvert\bm{X}}(t_{i}\rvert\bm{x}_{i})}\right]$
@@
In this expression, the first term represents the log-likelihood of observing the event at time $t_{i}$. The second term captures the conditional probability of observing the censoring time after the event time, given that the event occurred at $t_{i}$. The third and fourth terms, by symmetry, represent the same quantities for the censoring time. Given $C(u_{1},u_{2})$ is an Archimedean copula (e.g., Clayton or Frank), the partial derivatives in Equation F.4 have closed-form solutions, allowing the log-likelihood to be expressed in a closed form that can be optimized using gradient-based methods. To estimate F.4, we adopt the approach as described in Algorithm 1 in Foomani et al. [2023].
Copula selection. After training, for each model (marginals and copula), we choose the model that minimizes the Akaike Information Criterion (AIC) given the model’s negative log-likelihood (NLL) and its number of parameters. The AIC is a measure used to compare models, taking into account both the goodness of fit (via the likelihood) and model complexity (via the number of parameters). This offers a practical, data-driven approach to copula selection, as minimizing the NLL provides an objective criterion for which model best captures the dependency structure explicitly [Emura and Chen, 2018, Ch. 4]. We then extract $C_{\theta}$ from the chosen copula model and use that to calculate the dependent evaluation metrics.

### v3

@@ D.2 Integrated Brier Score
$\text{IBS}\left(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{S}_{T}(\cdot\mid\cdot),G(\cdot)\right)\ =\ \frac{1}{t_{\text{max}}}\cdot\int_{0}^{t_{\text{max}}}\text{BS}\left(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{S}_{T}(\cdot\mid\cdot),G(\cdot),\tau\right)d\tau,$ (D.4)
where $t_{\text{max}}$ is usually the maximum event time of the combined training and validation datasets. This metric is implemented in the SurvivalEVAL package [45].
D.3 Mean Absolute Error
@@
MAE-Margin [22] estimates a “best guess”-event time for censored subjects using the KM estimator [6]. For a subject censored at time $t_{i}$, the margin event time is computed as:
@@
This metric is implemented in the SurvivalEVAL package [45].
Appendix E Experimental details
E.1 Data generation and preprocessing
@@
Synthetic: This paper utilizes a linear synthetic dataset with two competing events (i.e., event and censoring), both based on the parametric Weibull distribution, as discussed in [17]. The datasets are characterized by the conditional hazard function and conditional survival probability of an event (or censoring) as follows:
$h_{T}(t\mid\bm{x}_{i})$ $=\left(\frac{v}{\rho}\right)\left(\frac{t}{\rho}\right)^{v-1}\exp\left(g_{\Psi}(\bm{x}_{i})\right),$ (E.1)
$S_{T}(t\mid\bm{x}_{i})$ $=\exp\left(-\left(\frac{t}{\rho}\right)^{v-1}\exp\left(g_{\Psi}(\bm{x}_{i})\right)\right),$ (E.2)
@@
where $v$ and $\rho$ represent the shape and scale parameters of the Weibull distribution, respectively. We use $v_{T}=4$, $v_{C}=6$, $\rho_{T}=17$, $\rho_{C}=19$. $g_{\Psi}(\bm{x})$ represents the risk function, which is a linear function in this work. The data generation process begins by initializing a copula with a predefined parameter $\theta$. We then sample $N$ sets of CDF observations from the copula, each containing two CDF probabilities (corresponding to each competing event, i.e., event or censoring). From these probabilities, we then calculate the corresponding times using the inverse function of (E.2). The data generation process is further described in Algorithm 2 in [17].
@@
METABRIC: The Molecular Taxonomy of Breast Cancer International Consortium (METABRIC) dataset [35] contains survival information for breast cancer patients. This dataset includes a diverse range of feature sets that encompass clinical traits, expression profiles, copy number variation (CNV) profiles, and single-nucleotide polymorphism (SNP) genotypes. All these features are derived from breast tumor samples collected during the METABRIC trial. The dataset can be downloaded from (https://www.cbioportal.org/study/summary?id=brca_metabric), and it does not have any missing values.
@@
MIMIC-IV: The Medical Information Mart for Intensive Care (MIMIC)-IV [36] contains critical care data from patients admitted to hospitals and intensive care units (ICU). We create a version using the MIMIC-IV database, which contains patients who are alive at least 24 hours after being admitted to ICU. Their date of death is derived from hospital records or state records, which means the cause of mortality is not limited to the reason for ICU admission. We follow the instructions in [23] to process the dataset, and the code is available in their GitHub repository: https://github.com/shi-ang/CensoredMAE.
@@
SEER: The SEER Program dataset [37] is a comprehensive collection of cancer patient data from approximately 49% of the U.S. population. Our study focuses on three subsets: SEER-brain, SEER-liver, and SEER-stomach, containing data on patients diagnosed with brain, liver, and stomach cancers, respectively. The goal is to model the time from diagnosis to failure events, such as death or disease progression. We preprocess the dataset as in [47]. Features analyzed include age, sex, cancer stage, grade, and various clinical factors. We excluded features with over 70% missing data, patients without follow-up times, duplicates, and those with a survival time of zero. The SEER cohort is available for download at www.seer.cancer.gov.
E.2 Survival learners
@@
CoxPH: The Cox Proportional Hazards model (CoxPH) is a semi-parametric survival analysis model [9], consisting of a non-parametric baseline hazard function $h_{0}(t)$ and a parametric partial hazard function $f(\bm{\theta},\bm{x}_{i})$, typically estimated using the Breslow method [48]. The hazard function is $h(t\mid\bm{x}_{i})=h_{0}(t)\exp(f(\bm{\theta},\bm{x}_{i}))$, where $f$ is linear. The MLE of $\hat{\bm{\theta}}$ maximizes the Cox partial log-likelihood [9]. CoxPH is implemented from the scikit-survival package version 0.21.0.
GBSA: GBSA [39] is a likelihood-based boosting model that optimizes an $\ell_{2}$-norm penalized Cox partial log-likelihood [9]. It updates candidate variables flexibly at each boosting step using an offset-based gradient boosting method, unlike traditional gradient boosting, which updates either a single component or all features for fitting the gradient. GBSA is implemented from the scikit-survival package version 0.21.0.
RSF: Random Survival Forests (RSF) [15] extend decision trees for survival analysis by constructing multiple trees on bootstrapped samples. Each tree predicts an individual’s survival function, with splits based on criteria like the log-rank test. RSF handles censored data as right-censored and accounts for it during tree construction, with final predictions as the average of all trees’ survival functions. RSF is implemented from the scikit-survival package version 0.21.0.
DeepSurv: DeepSurv [13] is a neural network implementation of the CoxPH model of the form $h\left(t\mid\bm{x}_{i}\right)=h_{0}\left(t\right)\exp\left(f\left(\bm{\theta},\bm{x}_{i}\right)\right)$, where $f\left(\bm{\theta},\bm{x}_{i}\right)$ denotes a risk score as a nonlinear function of the features, i.e.$f\left(\bm{\theta},\bm{x}_{i}\right)=\sigma(\bm{x}_{i}\bm{\theta})$, where $\sigma$ is a nonlinear activation function. Popular functions include the hyperbolic tangent (Tanh) function or the Rectified Linear Unit (ReLU). The MLE for $\hat{\bm{\theta}}$ is derived by numerically maximizing the partial Cox log-likelihood. DeepSurv is implemented from https://github.com/shi-ang/BNN-ISD/.
MTLR: Multi-Task Logistic Regression (MTLR) [16] is a neural network implementation of a discrete-time model that predicts the survival distribution using a sequence of dependent logistic regressors. The number of discrete times is determined by the square root of the number of uncensored patients, and we use quantiles to divide those uncensored instances evenly into each time interval. MTLR is implemented from https://github.com/shi-ang/BNN-ISD/.
@@
E.3 Implementation details
@@
Copula model. To validate our approach in semi-synthetic datasets, we train two copula models, Clayton and Frank, alongside two nonlinear Weibull proportional-hazards models for the marginal distributions $T$ and $C$, respectively. The Weibull models use a multilayer perceptron (MLP) as the backbone network architecture. We use a single layer with 32 hidden nodes and the ReLU activation function. For all experiments, we train the models for 10000 epochs using the Adam optimizer [49] with a learning rate of 1e-3, batch normalization enabled, and a batch size equal to the number of training instances. We use early stopping to terminate the training process if the validation loss does not improve for 100 consecutive epochs.
@@
Loss function. We estimate the model parameters using maximum likelihood estimation (MLE) under dependent censoring, as described in [17]. Given survival data $\mathcal{D}$, let $f_{T\mid\bm{x}}$ and $F_{T\mid\bm{x}}$ represent the conditional density and cumulative distribution functions, respectively, over the event horizon, with the event distribution $T$ and the censoring distribution ${C}$. Under independence censoring, the likelihood of the $i$-th instance experiencing the event is:
$\mathcal{L}(\mathcal{D})=\prod_{i=1}^{N}\left[\text{Pr}(T=t_{i},C\geq t_{i}\mid\bm{x}=\bm{x}_{i})^{\delta_{i}}\times\text{Pr}(T\geq t_{i},{C}=t_{i}\mid\bm{x}=\bm{x}_{i})^{1-\delta_{i}}\right].$ (E.3)
@@
However, when $T$ and $C$ are no longer independent, we can no longer rely on this clean decomposition of the survival likelihood. Based on Sklar’s theorem [42], in order to adjust Equation E.3 to dependent censoring under an Archimedean copula, we introduce the following lemma [17]:
Lemma 1.
Under Sklar’s (Survival) Theorem, and given $C(u_{1},u_{2})$ is a Archimedean copula, then if
$S_{T,C}(t,c|\bm{x})=C(u_{1},u_{2})\Big{|}^{u_{1}=S_{T}(t\rvert\bm{x})}_{u_{2}=S_{C}(c\rvert\bm{x})},$ (E.4)
then we have
$\int_{c}^{\infty}f_{C}(c\mid t,\bm{x})=\frac{\partial}{\partial u_{1}}C(u_{1},u_{2})\Big{|}^{u_{1}=S_{T}(t\rvert\bm{x})}_{u_{2}=S_{C}(c\rvert\bm{x})}.$ (E.5)
Applying Lemma 1 to Equation E.3 and taking the sum of logs yields the log-likelihood for survival models under dependent censoring:
@@
$\mathcal{L}(\mathcal{D})=\sum_{i=1}^{N}\delta_{i}\;\text{log}\left[f_{T}(t_{i}|\bm{x}_{i})\right]\;+\;\delta_{i}\;\text{log}\left[\frac{\partial}{\partial u_{1}}C(u_{1},u_{2})\rvert^{u_{1}=S_{T}(t_{i}\rvert\bm{x}_{i})}_{u_{2}=S_{C}(t_{i}\rvert\bm{x}_{i})}\right]$
$+\;(1-\delta_{i})\;\text{log}\left[f_{C}(t_{i}|\bm{x}_{i})\right]+(1-\delta_{i})\;\text{log}\left[\frac{\partial}{\partial u_{2}}C(u_{1},u_{2})\rvert^{u_{1}=S_{T\rvert}(t_{i}\rvert\bm{x}_{i})}_{u_{2}=S_{C}(t_{i}\rvert\bm{x}_{i})}\right].$ (E.6)
@@
In this expression, the first term represents the log-likelihood of observing the event at time $t_{i}$. The second term captures the conditional probability of observing the censoring time after the event time, given that the event occurred at $t_{i}$. The third and fourth terms, by symmetry, represent the same quantities for the censoring time. Given $C(u_{1},u_{2})$ is an Archimedean copula (e.g., Clayton or Frank), the partial derivatives in Equation E.3 have closed-form solutions, allowing the log-likelihood to be expressed in a closed form that can be optimized using gradient-based methods. To estimate Equation E.3, we follow the approach described by Foomani et al. [17], as outlined in Algorithm 1. The parameters are optimized using gradient descent with the Adam optimizer [49].
Algorithm 1 Learning Under Dependent Censoring. Reproduced from [17].
1:Dataset $\mathcal{D}=\left\{\left(\bm{x}_{i},t_{i},\delta_{i}\right)\right\}_{i=1}^{N}$; Copula $C_{\theta}$: bivariate copula parameterized by $\theta$; $M$: survival model class parameterized by $\phi$ that can produce $\hat{S}^{(M)}_{T}(t|\bm{x})$ and $\hat{f}^{(M)}_{T}(t|\bm{x})$; $\alpha$: learning rate; $M$: number of training epochs; $K$: large constant; $\theta_{\min}$: small positive threshold for $\theta$.
2:$\hat{\theta},\hat{\phi}_{T},\hat{\phi}_{C}$: learned parameters of the copula and each marginal survival model.
3:$M_{T}\leftarrow\text{Instantiate}(M;\hat{\psi}^{(0)}_{T})$
4:$M_{C}\leftarrow\text{Instantiate}(M;\hat{\psi}^{(0)}_{C})$
5:$C_{\theta}\leftarrow\text{Instantiate}(C;\hat{\theta}^{(0)})$
6:for $i=1$ to $M$ do
7:     $L_{i}\leftarrow\ell(D;\hat{f}^{(M_{T})}_{T},\hat{f}^{(M_{C})}_{T},\hat{S}^{(M_{T})}_{T},\hat{S}^{(M_{C})}_{T},C_{\hat{\theta}^{(i)}})$
8:     $\hat{\psi}_{C}^{(i)}\leftarrow\text{Adam}(L_{i},\hat{\psi}_{C},\alpha)$
9:     $\hat{\psi}_{T}^{(i)}\leftarrow\text{Adam}(L_{i},\hat{\psi}_{T},\alpha)$
10:     $\nabla\hat{\theta}^{(i)}\leftarrow\nabla\hat{\theta}^{(i)}\times K$
11:     $\nabla\hat{\theta}^{(i)}\leftarrow\nabla\hat{\theta}^{(i)}\big{|}_{[-0.1,0.1]}$
12:     $\hat{\theta}^{(i)}\leftarrow\text{Adam}(L_{i},\hat{\theta},\alpha)$
13:     $\hat{\theta}^{(i)}\leftarrow\min(\hat{\theta}^{(i)},\theta_{\min})$
14:end for
15:return $\hat{\theta}^{(i)},\hat{\psi}_{T}^{(i)},\hat{\psi}_{C}^{(i)}$